# A Bombieri-Vinogradov theorem for sectors in real quadratic number fields

### v1

@@ 6. Estimation of type II sums
(6.1) $\begin{split}T_{II}(M)\ll&\Big{(}M+P^{8/3}\tilde{N}M^{1/3}\log^{2}2P\Big{)}^{1/2}\Big{(}xM^{-1}+P^{8/3}\tilde{N}x^{1/3}M^{-1/3}\log^{2}2P\Big{)}^{1/2}\times\\ &\Big{(}\sum\limits_{\begin{subarray}{c}\mathcal{N}(\mathfrak{m})\leq 2M\end{subarray}}|b_{\mathfrak{m}}|^{2}\Big{)}^{1/2}\Big{(}\sum\limits_{\begin{subarray}{c}\mathcal{N}(\mathfrak{r})\leq x/M\end{subarray}}|c_{\mathfrak{r}}|^{2}\Big{)}^{1/2}\log^{2}x.\end{split}$
@@
For any $K,L\geq 1$ and sequences $b_{\mathfrak{m}}$ and $c_{\mathfrak{r}}$ of complex numbers with indices $\mathfrak{m}$ and $\mathfrak{r}$ running over $\mathcal{I}_{\mathbb{K}}$, we have, using Cauchy’s inequality and Lemma 6,
(6.2) $\begin{split}&\sum\limits_{\tilde{\chi}\in F(P,N)}\Big{|}\sum\limits_{\begin{subarray}{c}\mathcal{N}(\mathfrak{m})\leq K\end{subarray}}\sum\limits_{\begin{subarray}{c}\mathcal{N}(\mathfrak{r})\leq L\end{subarray}}b_{\mathfrak{m}}c_{\mathfrak{r}}\tilde{\chi}(\mathfrak{m}\mathfrak{r})\Big{|}\\ \leq&\Big{(}\sum\limits_{\tilde{\chi}\in F(P,N)}\Big{|}\sum\limits_{\begin{subarray}{c}\mathcal{N}(\mathfrak{m})\leq K\end{subarray}}b_{\mathfrak{m}}\tilde{\chi}(\mathfrak{m})\Big{|}^{2}\Big{)}^{1/2}\Big{(}\sum\limits_{\tilde{\chi}\in F(P,N)}\Big{|}\sum\limits_{\begin{subarray}{c}1\leq\mathcal{N}(\mathfrak{r})\leq L\end{subarray}}c_{\mathfrak{r}}\tilde{\chi}(\mathfrak{r})\Big{|}^{2}\Big{)}^{1/2}\\ \ \ll&\left(K+P^{8/3}\tilde{N}K^{1/3}\log^{2}2P\right)^{1/2}\left(L+P^{8/3}\tilde{N}L^{1/3}\log^{2}2P\right)^{1/2}\times\\ &\Big{(}\sum\limits_{\begin{subarray}{c}\mathcal{N}(\mathfrak{m})\leq K\end{subarray}}|b_{\mathfrak{m}}|^{2}\Big{)}^{1/2}\Big{(}\sum\limits_{\begin{subarray}{c}\mathcal{N}(\mathfrak{r})\leq L\end{subarray}}|c_{\mathfrak{r}}|^{2}\Big{)}^{1/2}.\end{split}$
The term $T_{II}(M)$ defined in (4.7) takes the same form as the sum in the first line above with $K=2M$ and $L=x/M$, except that there are two extra summation conditions $\mathcal{N}(\mathfrak{m})\leq y/V$ and $U<\mathcal{N}(\mathfrak{mr})\leq y$ and an extra maximum over $y\leq x$. These extra conditions and maximum can be removed in a standard way using Proposition 4 (cosmetic surgery), reducing the estimation of $T_{II}(M)$ to (6.2) (for details, see [3, proof of Satz 5.5.2.], for example). This comes at the cost of an extra factor of $\log^{2}x$, as compared to the bound in (6.2). Hence, we arrive at the desired estimate (6.1). ∎
@@
$\sum\limits_{\begin{subarray}{c}\mathcal{N}(\mathfrak{m})\leq K\end{subarray}}|b_{\mathfrak{m}}|^{2}=\sum\limits_{\begin{subarray}{c}\mathcal{N}(\mathfrak{m})\leq K\end{subarray}}\Lambda({\mathfrak{m}})^{2}\ll K\log 2K$
@@
$\sum\limits_{\mathcal{N}(\mathfrak{r})\leq L}|c_{\mathfrak{r}}|^{2}=\sum\limits_{\mathcal{N}(\mathfrak{r})\leq L}\tau(\mathfrak{r})^{2}\ll L\cdot\log^{3}2L,$
where $\tau(\mathfrak{r})$ denotes the number of prime ideal divisors of $\mathfrak{r}$, we deduce that
(6.3) $\begin{split}T_{II}(M)\ll&\left(M^{1/2}+P^{4/3}\tilde{N}^{1/2}M^{1/6}\log 2P\right)\left(x^{1/2}M^{-1/2}+P^{4/3}\tilde{N}^{1/2}x^{1/6}M^{-1/6}\log 2P\right)\times\\ &(M\log x)^{1/2}(xM^{-1}\log^{3}x)^{1/2}\log^{2}x\\ \ll&\left(x+x^{2/3}M^{1/3}P^{4/3}\tilde{N}^{1/2}\log 2P+xM^{-1/3}P^{4/3}\tilde{N}^{1/2}\log 2P+x^{2/3}P^{8/3}\tilde{N}\log^{2}2P\right)\log^{4}x.\end{split}$
@@
(6.4) $T_{II}\ll\Big{(}x+xP^{4/3}\tilde{N}^{1/2}(U^{-1/3}+V^{-1/3})+x^{2/3}P^{8/3}\tilde{N}\Big{)}\log^{7}x,$
@@ 7. Proof of Theorem 1
Combining (3.35), (4.2), (6.4) and (4.5) we obtain
(7.1) $D(P,N)\ll\Big{(}xP^{-1}\tilde{N}^{-1}+xP^{1/3}\tilde{N}^{-1/2}(U^{-1/3}+V^{-1/3})+x^{2/3}P^{5/3}+x^{1/3}P^{4/3}(UV)^{2/3}+PU\Big{)}\log^{9}x$
if $P\leq x$. Now we assume, more restrictively, that $P\leq x^{2/3}$ and pick
$U:=V:=x^{2/5}P^{-3/5},$
noting that $U,V\geq 1$ under this assumption $P\leq x^{2/3}$. Under this choice, (7.1) implies
$D(P,N)\ll\left(xP^{-1}+x^{13/15}P^{8/15}+x^{2/3}P^{5/3}\right)\log^{9}x.$
@@
$\sum\limits_{2\leq\mathcal{N}(\mathfrak{q})\leq Q}E^{\ast}(x;\mathfrak{q})\ll_{A}\left(x^{13/15}Q^{8/15}+x^{2/3}Q^{5/3}\right)\log^{11}x+\frac{x}{\log^{A}x}$
if $Q\leq x^{2/3}$. Recalling the notations in subsection 3.1, this is the same as the bound
(7.2) $\sum\limits_{\begin{subarray}{c}\mathfrak{q}\in\mathcal{I}_{\mathbb{K}}\\ \mathcal{N}(\mathfrak{q})\leq Q\end{subarray}}\,\,\max\limits_{\begin{subarray}{c}a\bmod\mathfrak{q}\\ (a,\mathfrak{q})=1\end{subarray}}\,\,\max\limits_{y\leq x}\left|\tilde{E}_{S}(y;\mathfrak{q},a)\right|\ll_{A}\left(x^{13/15}Q^{8/15}+x^{2/3}Q^{5/3}\right)\log^{11}x+\frac{x}{\log^{A}x}.$
Thus, for the proof of Theorem 1, it suffices to approximate $\tilde{E}_{S}(y;\mathfrak{q},a)$ sufficiently closely by $E_{S}(y;\mathfrak{q},a)$. To this end, we need to appoximate the main term
@@
$M_{S}(y;\mathfrak{q},a)=\tilde{M}_{S}(y;\mathfrak{q},a)+O\left(\frac{x}{\log^{A+1}x}\right)$
@@
Remark 4. The sole reason that we are not able to handle small sectors with $\eta=x^{-\delta}$ for $\delta$ in a suitable interval $0<\delta\leq\delta_{0}$ along the lines above is the following. Our treatment of small moduli with $\mathcal{N}(\mathfrak{q})\leq Q_{0}=\log^{A}x$ making direct use of the Siegel-Walfisz theorem does not give a bound better than $O(x\exp(-C\sqrt{\log x}))$ for their total contribution, no matter what the size of $\eta$ is. However, we need a bound of order of magnitude $O(\eta x\log^{-A}x)$ if $\eta$ is allowed to depend on $x$. To resolve this issue, we may resort to using the explicit formula and zero-density estimates for Hecke $L$-functions, which requires more machinery. Of course, large sieve inequalities for Hecke characters (more generally, mean value estimates for Dirichlet polynomials twisted with Hecke characters) are needed here as well.
Acknowledgements. The first-named author would like to thank the Ramakrishna Mission Vivekananda Educational and Research Institute for an excellent work environment. The second-named author is grateful to the Indian Statistical Institute Kolkata for its excellent research environment and acknowledges support from the Research Associate Fellowship at the Institute.

### v2

@@ 6. Estimation of type II sums
(6.1) $\begin{split}T_{II}(M)\ll&\Big(M+P^{3}\tilde{N}(Px)^{\varepsilon}\Big)^{1/2}\Big(xM^{-1}+P^{3}\tilde{N}(Px)^{\varepsilon}\Big)^{1/2}\times\\ &\Big(\sum\limits_{\begin{subarray}{c}\mathcal{N}(\mathfrak{m})\leq 2M\end{subarray}}|b_{\mathfrak{m}}|^{2}\Big)^{1/2}\Big(\sum\limits_{\begin{subarray}{c}\mathcal{N}(\mathfrak{r})\leq x/M\end{subarray}}|c_{\mathfrak{r}}|^{2}\Big)^{1/2}\log^{2}x.\end{split}$
@@
For any $K,L\geq 1$ and sequences $b_{\mathfrak{m}}$ and $c_{\mathfrak{r}}$ of complex numbers with indices $\mathfrak{m}$ and $\mathfrak{r}$ running over $\mathcal{I}_{\mathbb{K}}$, we have, using the Cauchy-Schwarz inequality and Lemma 6,
(6.2) $\begin{split}&\sum\limits_{\tilde{\chi}\in F(P,N)}\Big|\sum\limits_{\begin{subarray}{c}\mathcal{N}(\mathfrak{m})\leq K\end{subarray}}\sum\limits_{\begin{subarray}{c}\mathcal{N}(\mathfrak{r})\leq L\end{subarray}}b_{\mathfrak{m}}c_{\mathfrak{r}}\tilde{\chi}(\mathfrak{m}\mathfrak{r})\Big|\\ \leq&\Big(\sum\limits_{\tilde{\chi}\in F(P,N)}\Big|\sum\limits_{\begin{subarray}{c}\mathcal{N}(\mathfrak{m})\leq K\end{subarray}}b_{\mathfrak{m}}\tilde{\chi}(\mathfrak{m})\Big|^{2}\Big)^{1/2}\Big(\sum\limits_{\tilde{\chi}\in F(P,N)}\Big|\sum\limits_{\begin{subarray}{c}1\leq\mathcal{N}(\mathfrak{r})\leq L\end{subarray}}c_{\mathfrak{r}}\tilde{\chi}(\mathfrak{r})\Big|^{2}\Big)^{1/2}\\ \ \ll&\left(K+P^{3}\tilde{N}(PK)^{\varepsilon}\right)^{1/2}\left(L+P^{3}\tilde{N}(PL)^{\varepsilon}\right)^{1/2}\times\\ &\Big(\sum\limits_{\begin{subarray}{c}\mathcal{N}(\mathfrak{m})\leq K\end{subarray}}|b_{\mathfrak{m}}|^{2}\Big)^{1/2}\Big(\sum\limits_{\begin{subarray}{c}\mathcal{N}(\mathfrak{r})\leq L\end{subarray}}|c_{\mathfrak{r}}|^{2}\Big)^{1/2}.\end{split}$
The term $T_{II}(M)$ defined in (4.7) takes the same form as the sum in the first line above with $K=2M$ and $L=x/M$, except that there are two extra summation conditions $\mathcal{N}(\mathfrak{m})\leq y/V$ and $U<\mathcal{N}(\mathfrak{mr})\leq y$ and an extra maximum over $y\leq x$. These extra conditions and maximum can be removed in a standard way using Lemma 4 (cosmetic surgery), reducing the estimation of $T_{II}(M)$ to (6.2) (for details, see [3, proof of Satz 5.5.2.], for example). This comes at the cost of an extra factor of $\log^{2}x$, as compared to the bound in (6.2). Hence, we arrive at the desired estimate (6.1). ∎
@@
$\sum\limits_{\begin{subarray}{c}\mathcal{N}(\mathfrak{m})\leq K\end{subarray}}|b_{\mathfrak{m}}|^{2}\leq\sum\limits_{\begin{subarray}{c}\mathcal{N}(\mathfrak{m})\leq K\end{subarray}}\Lambda({\mathfrak{m}})^{2}\ll K\log 2K$
@@
$\sum\limits_{\mathcal{N}(\mathfrak{r})\leq L}|c_{\mathfrak{r}}|^{2}\leq\sum\limits_{\mathcal{N}(\mathfrak{r})\leq L}\tau(\mathfrak{r})^{2}\ll L\cdot\log^{3}2L,$
where $\tau(\mathfrak{r})$ denotes the number of ideal divisors of $\mathfrak{r}$, we deduce that
(6.3) $\begin{split}T_{II}(M)\ll&\left(M^{1/2}+P^{3/2}\tilde{N}^{1/2}(Px)^{\varepsilon/2}\right)\left(x^{1/2}M^{-1/2}+P^{3/2}\tilde{N}^{1/2}(Px)^{\varepsilon/2}\right)\times\\ &(M\log x)^{1/2}(xM^{-1}\log^{3}x)^{1/2}\log^{2}x\\ \ll&\Big(x+xP^{3/2}\tilde{N}^{1/2}M^{-1/2}(Px)^{\varepsilon/2}+x^{1/2}P^{3/2}\tilde{N}^{1/2}M^{1/2}(Px)^{\varepsilon/2}\\ &+x^{1/2}P^{3}\tilde{N}(Px)^{\varepsilon}\Big)\log^{4}x.\end{split}$
@@
(6.4) $T_{II}\ll\Big(x+xP^{3/2}\tilde{N}^{1/2}\left(U^{-1/2}+V^{-1/2}\right)(Px)^{\varepsilon/2}+x^{1/2}P^{3}\tilde{N}(Px)^{\varepsilon}\Big)\log^{7}x,$
@@ 7. Proof of Theorem 1
Combining (3.36), (4.2), (4.5) and (6.4), we obtain
(7.1) $\begin{split}D(P,N)\ll&\Big(xP^{-1}\tilde{N}^{-1}+xP^{1/2}\tilde{N}^{-1/2}(U^{-1/2}+V^{-1/2})(Px)^{\varepsilon/2}+\\ &x^{1/2}P^{2}(Px)^{\varepsilon}+x^{1/3}P^{4/3}(UV)^{2/3}+PU\Big)\log^{9}x\end{split}$
if $P\leq x$. Now we assume, more restrictively, that $P\leq x^{4/5}$ and pick
$U:=V:=x^{4/11}P^{-5/11},$
noting that $U,V\geq 1$ under this assumption. Under this choice, (7.1) implies
$D(P,N)\ll\left(xP^{-1}+\left(x^{9/11}P^{8/11}+x^{1/2}P^{2}\right)(Px)^{\varepsilon}\right)\log^{9}x.$
@@
$\sum\limits_{2\leq\mathcal{N}(\mathfrak{q})\leq Q}E^{\ast}(x;\mathfrak{q})\ll\left(x^{9/11}Q^{8/11}+x^{1/2}Q^{2}\right)x^{2\varepsilon}+\frac{x}{\log^{A}x}$
if $Q\leq x^{4/5}$. Recalling the notations in subsection 3.1, this is the same as the bound
(7.2) $\sum\limits_{\mathcal{N}(\mathfrak{q})\leq Q}\,\,\max\limits_{\begin{subarray}{c}a\bmod\mathfrak{q}\\ (a,\mathfrak{q})=1\end{subarray}}\,\,\max\limits_{y\leq x}\left|\tilde{E}_{S}(y;\mathfrak{q},a)\right|\ll\left(x^{9/11}Q^{8/11}+x^{1/2}Q^{2}\right)x^{\varepsilon}+\frac{x}{\log^{A}x}$
upon redefining $\varepsilon$. Thus, for the proof of Theorem 1, it suffices to approximate $\tilde{E}_{S}(y;\mathfrak{q},a)$ sufficiently closely by $E_{S}(y;\mathfrak{q},a)$. To this end, we need to appoximate the main term
@@
$M_{S}(y;\mathfrak{q},a)=\tilde{M}_{S}(y;\mathfrak{q},a)+O\left(\frac{x}{\varphi(\mathfrak{q})\log^{A+1}x}\right)$
@@
Remark 4. The sole reason that we are not able to handle small sectors with $\eta=x^{-\delta}$ for $\delta$ in a suitable interval $0<\delta\leq\delta_{0}$ along the above lines is the following. Our treatment of small moduli with $\mathcal{N}(\mathfrak{q})\leq Q_{0}=\log^{A}x$ making direct use of the Siegel-Walfisz theorem does not give a bound better than $O(x\exp(-C\sqrt{\log x}))$ for their total contribution, no matter what the size of $\eta$ is. However, we need a bound of order of magnitude $O(\eta x\log^{-A}x)$ if $\eta$ is allowed to depend on $x$. To resolve this issue, we may resort to using the explicit formula and zero-density estimates for Hecke $L$-functions, which requires more machinery. Of course, large sieve inequalities for Hecke characters (more generally, mean value estimates for Dirichlet polynomials twisted with Hecke characters) are needed here as well.
Acknowledgements. Both authors would like to thank the anonymous referee for valuable comments, in particular, for pointing out a way to improve our initial level of distribution $1/5-\varepsilon$ to $1/4-\varepsilon$ using a refined treatment of the large sieve with Hecke characters. The first-named author would like to thank the Ramakrishna Mission Vivekananda Educational and Research Institute for an excellent work environment. The second-named author is grateful to the Indian Statistical Institute Kolkata for its excellent research environment and acknowledges support from the Research Associate Fellowship at the Institute.